# Compact and stable source of polarization-entangled photon-pairs based on a folded linear displacement interferometer

### v1

@@
The realization of quantum networks requires the development of robust low size, weight and power (SWaP) systems suitable for operation under harsh environments in remote and mobile nodes such as satellites. We present a source of polarization-entangled photon-pairs in a folded linear displacement interferometer based on spontaneous parametric down conversion using a Type-0 periodically poled potassium titanyl phosphate crystal. Featuring a compact and stable double-pass geometry using a corner-cube retroreflector, the source has a detected pair rate of 2.5 M pairs/s/mW with a Bell state fidelity of 94.1% $\pm$ 2.1%. The quality and demonstrated performance of the source are suitable for deployment in entanglement-based quantum networks.
@@ 3 Experiment
The experimental setup for the FLDI source (Fig. 2(a)) is divided into four sections, with each section performing a different role within the source. They are (1) the pump power and polarization control, (2) the beam steering, (3) entanglement generation and, finally, (4) the detection. A single-mode, broadband continuous-wave laser at 405 nm with a spectral bandwidth of 0.7 nm (Thorlabs LP405C1) pumps a Type-0 PPKTP (Raicol Crystals) of dimensions $1\times 2\times 10$ mm and poling-period, $\Lambda$ = 3.425 $\mu$m. A fluorescent filter (FF) is used to filter unwanted emission from the pump. A combination of a half-wave plate (HWP1) and a polarizing beam splitter (PBS) is used for optical control of pump power. A second half-wave plate (HWP2) is used to adjust the polarization of the beam. A liquid crystal variable retarder (LCVR) controls the phase between the two photon-pairs generated by adjusting the H and V components of the pump beam.
The pump light is focused into the PPKTP by a planoconvex lens (pump focussing lens, PFL) of focal length 200 mm. The beam steering consists of two mirrors that direct the pump into the PPKTP. A beam displacer (BD) crystal ($\alpha$-BBO, AG Optics) of length 13 mm is used to split the incoming beam into two parallel beams with orthogonal polarizations to each other separated by approximately 1 mm. The orthogonal polarization components then follow the entanglement generation scheme described above. This results in a compact interferometer with an overall length of $\sim 9.5$ cm (Fig. 2 (b)) whilst still allowing access for measurement of the power incident on the PPKTP. Thus, there is scope for further size reduction with optimized mounting of the components, for example, in a packaged system.
@@
Figure 2: FLDI configuration.(a) Experimental schematic of the FLDI source in 4 sections: (1) pump preparation, (2) beam steering, (3) entanglement generation and (4) detection. FF: Fluorescent filter, HWP: half-wave plate, PBS: polarizing beam splitter, PFL: pump focusing lens, DM: dichroic mirror, BD: beam displacer, CCR: cornercube retroreflector, LPF: long pass filter, BPF: bandpass filter, P: linear polarizer, D: Single photon detectors, CC: coincidence counter. (b) Photograph of section (3), the entanglement generation. Featuring DM1, BD1, a-HWP, PPKTP and CCR, the interferometer has a of length $\sim 9.5$ cm.
@@ 4 Results
Multiple peaks are observed in both contour plots; Due to the trade-off between the pair rate and the heralding efficiency in the SPDC process, the spectral conditions for optimal pair rate do not necessarily overlap with those of the heralding efficiency. The focusing of the pump and SPDC modes in the crystal is quantified by a dimensionless parameter $\xi$, defined as [29]
@@
where $L$ is the length of the crystal, $k_{j}$ and $w_{j}$ are the respective wavevectors and beam waists of pump, signal and idler modes. Change in the spectra of pump and SPDC varies the focal parameter $\xi$, showing visible peaks and troughs in the contour plots. For a given fixed waists for all the three modes, Stronger focusing of the pump in the SPDC crystal enhances mode overlap leading to an optimal pair production with reduced heralding efficiencies. For a chosen optimal condition from (a) and (b), the measured spectra of the SPDC photons and the pump beam are shown in Figs. 2(c) and (d). The laser was heated to $40^{\circ}$ C, corresponding to a central wavelength of 404.6 nm. The SPDC spectrum is broadband with multiple peaks observed near degenerate (810 nm) and near non-degenerate regimes. The largest peak is the signal at 781.3 nm, with smaller peak for the idler at 845.6 nm. Placing a 10 nm bandpass filter centred at 780 nm ensures detection of signal photon at 780 nm and corresponding phase matched idler photon.
@@
To quantify the entanglement produced by the source, polarization correlation measurements on photon-pairs were carried out using a calibrated polarizer on both the signal (P1) and idler (P2) arms. In this two-linear polarizer setting, the signal photon is projected into one polarization state using a fixed polarizer (P1) and the polarizer on the idler detector arm is swept through 180 degrees. With the angle P1 set to zero (H projection), the coincidences were recorded for different angles P2. Similar measurements were carried out for P1 angles 45, 90 & 135, which correspond to D, V & A polarization projections. Fig. 5 shows the plots of variation of the coincidences with the angle P2 for the projections H, V, D & A projections on P1. From the plots, the visibilities were estimated to be 98.8%$\pm$1.2(H), 97.0%$\pm$1.1(V), 89.3%$\pm$1.0(D) & 91.2%$\pm$1.0 %(A) and the fidelity with $|\Phi^{+}\rangle$ is 94.1%$\pm$2.1%. With this, the quantum bit error rate is estimated to be 5.9%.
@@ 5 Discussion and Conclusion
Funding Innovate UK project NextSTEPS 10031944, EPSRC Research Excellence Award Studentship (EP/T517938/1), INSQT, EPSRC Quantum Technology Hub in Quantum Communication (EP/T001011/1), EPSRC International Network in Space Quantum Technologies (EP/W027011/1), and the EPSRC Integrated Quantum Networks Hub (EP/Z533208/1).

### v2

@@
The realization of quantum networks requires the development of robust low size, weight and power (SWaP) systems suitable for operation under harsh environments in remote and mobile nodes such as satellites. We present a source of polarization-entangled photon-pairs in a folded linear displacement interferometer based on spontaneous parametric down conversion using a Type-0 periodically poled potassium titanyl phosphate crystal. Featuring a compact and stable double-pass geometry using a corner-cube retroreflector, the source has a detected pair rate of 2.5 M pairs/s/mW with a Bell state fidelity of 94.1% $\pm$ 2.1%. The qualities and demonstrated performance of the source make it suitable for deployment in entanglement-based quantum networks.
@@ 3 Experiment
The experimental setup for the FLDI source (Fig. 2 (a)) is divided into four sections: (1) pump power and polarization control, (2) beam steering, (3) entanglement generation and, finally, (4) detection. A single-mode, broadband continuous-wave laser at 405 nm with a spectral bandwidth of 0.7 nm (Thorlabs LP405C1) pumps a Type-0 PPKTP (Raicol Crystals) of dimensions $1\times 2\times 10$ mm and poling-period, $\Lambda$ = 3.425 $\mu$m. A fluorescent filter (FF) is used to filter unwanted emission from the pump. A combination of a half-wave plate (HWP1) and a polarizing beam splitter (PBS) is used for optical control of pump power. A second half-wave plate (HWP2) is used to adjust the polarization of the beam. A liquid crystal variable retarder (LCVR) controls the phase between the two photon-pairs generated by adjusting the H and V components of the pump beam.
The pump light is focused into the PPKTP by a planoconvex lens (pump focussing lens, PFL) of focal length 200 mm. The beam steering consists of two mirrors that direct the pump into the PPKTP. A beam displacer (BD) crystal ($\alpha$-BBO, AG Optics) of length 13 mm is used to split the incoming beam into two parallel beams with orthogonal polarizations to each other separated by approximately 1 mm. The orthogonal polarization components then follow the entanglement generation scheme described in Section 2. This results in a compact interferometer with an overall length of $\sim 9.5$ cm (Fig. 2 (b)) whilst still allowing access for measurement of the power incident on the PPKTP. Thus, there is scope for further size reduction with optimized mounting of the components, for example, in a packaged system.
@@
Figure 2: FLDI configuration.(a) Experimental schematic of the FLDI source in four sections: (1) pump preparation, (2) beam steering, (3) entanglement generation and (4) detection. FF: Fluorescent filter, HWP: half-wave plate, PBS: polarizing beam splitter, PFL: pump focusing lens, DM: dichroic mirror, BD: beam displacer, CCR: cornercube retroreflector, LPF: long pass filter, BPF: bandpass filter, P: linear polarizer, D: Single photon detectors, CC: coincidence counter. (b) Photograph of section (3), the entanglement generation. Featuring DM1, BD1, a-HWP, PPKTP and CCR, the interferometer has a of length $\sim 9.5$ cm.
@@ 4 Results
Multiple peaks are observed in both contour plots. Due to the trade-off between the pair rate and the heralding efficiency in the SPDC process, the spectral conditions for optimal pair rate do not necessarily overlap with those of the heralding efficiency. The focusing of the pump and SPDC modes in the crystal is quantified by a dimensionless parameter $\xi$, defined as [29]
@@
where $L$ is the length of the crystal, $k_{j}$ and $w_{j}$ are the respective wavevectors and beam waists of pump, signal and idler modes. Change in the spectra of pump and SPDC affects the focal parameter $\xi$, showing visible peaks and troughs in the contour plots. For a given fixed waists for all the three modes, Stronger focusing of the pump in the SPDC crystal enhances mode overlap leading to an optimal pair production with reduced heralding efficiencies. For a chosen optimal condition from (a) and (b), the measured spectra of the SPDC photons and the pump beam are shown in Fig. 2 (c) and (d). The laser was heated to $40^{\circ}$ C, corresponding to a central wavelength of 404.6 nm. The SPDC spectrum is broadband with multiple peaks observed near degenerate (810 nm) and non-degenerate regimes. The largest peak is the signal at 781.3 nm, with smaller peak for the idler at 845.6 nm. Placing a 10 nm bandpass filter centred at 780 nm ensures detection of signal photon at 780 nm and corresponding phase matched idler photon.
@@
To quantify the entanglement produced by the source, polarization correlation measurements on photon-pairs were carried out using a calibrated polarizer on both the signal (P1) and idler (P2) arms. In this two-linear polarizer setting, the signal photon is projected into one polarization state using a fixed polarizer (P1) and the polarizer on the idler detector arm is swept through 180 degrees. With the angle P1 set to zero (H projection), the coincidences were recorded for different angles P2. Similar measurements were carried out for P1 angles 45, 90 & 135, which correspond to D, V & A polarization projections. Fig. 5 shows the plots of variation of the coincidences with the angle P2 for the projections H, V, D & A projections on P1. From the plots, the visibilities were estimated to be 98.8% $\pm$ 1.2% (H), 97.0% $\pm$ 1.1% (V), 89.3% $\pm$ 1.0% (D) & 91.2% $\pm$ 1.0% (A) and the fidelity with $|\Phi^{+}\rangle$ is 94.1% $\pm$ 2.1%. With this, the quantum bit error rate is estimated to be 5.9%.
@@ 5 Discussion and Conclusion
Funding Innovate UK project NextSTEPS 10031944, EPSRC Research Excellence Award Studentship (EP/T517938/1), EPSRC Quantum Technology Hub in Quantum Communication (EP/T001011/1), EPSRC International Network in Space Quantum Technologies (EP/W027011/1), and the EPSRC Integrated Quantum Networks Hub (EP/Z533208/1).